# Cert-SSB: Toward Certified Sample-Specific Backdoor Defense

### v1

@@ V Experiments
TABLE III: Certified performance ($i.e.$, ERA and AER) is reported using the best results across different noise levels at each radius under the all-to-all setting. We evaluate three types of attacks (one-pixel, four-pixel, and blending) on the MNIST, CIFAR-10, and ImageNette datasets. In particular, we mark the best certification results in boldface.
@@
Cert-SSB 0.70 93.32 78.16 42.48 52.34 38.26 17.28 1.40 0
TABLE IV: Certified performance ($i.e.$, CRA and ACR) is reported using the best results across different noise levels at each radius under the all-to-all setting. We evaluate three types of attacks (one-pixel, four-pixel, and blending) on the MNIST, CIFAR-10, and ImageNette datasets. In particular, we mark the best certification results in boldface.
@@ V-A2 Training Settings
We adopt a sample-specific smoothing approach during training. In this stage, we set the number of sampled Gaussian noise vectors ($i.e.$, augmented datasets) to $M=1,000$ for MNIST and CIFAR-10, and $M=200$ for ImageNette, resulting in ensembles of 1,000 and 200 models, respectively. Following previous works [59, 27], the added noise follows a Gaussian distribution with mean $\mu=0$ and standard deviation $\sigma$, set as follows: for MNIST and CIFAR-10, $\sigma\in\{0.12,0.25,0.5,1.0\}$; for ImageNette, $\sigma\in\{0.25,0.5,1.0\}$. Additionally, we set the number of stochastic gradient ascent iterations to $T=100$, the number of Monte Carlo samples to $J=1$, and the learning rate to $\alpha=10^{-4}$. We initialize $\sigma_{0}$ using the optimal sample-specific noise level $\sigma_{\bm{x}}^{*}$ obtained during training.
@@ V-A3 Attack Settings
We evaluate the certified performance of Cert-SSB against three representative backdoor attacks: one-pixel pattern, four-pixel pattern, and random but fixed noise patterns blended across the entire image [71]. The perturbation magnitude of the attack is controlled by the $\ell_{2}$-norm of the backdoor patterns, with $\left\|\Delta\right\|_{2}=0.1$. Following prior work [24], we inject 10% poisoned samples into the MNIST dataset and 5% into the CIFAR-10 and ImageNette datasets. The goal of these attacks is to induce the model to misclassify inputs as ‘0’ in MNIST, ‘airplane’ in CIFAR-10, and ‘tench’ in ImageNette. In addition to the all-to-one attack, we also consider an all-to-all attack objective [32], where the compromised model alters its predictions based on the original labels. We hereby primarily focus on the perturbation magnitude and the number of injected backdoor samples without considering specific backdoor patterns.
@@ V-A4 Evaluation Metrics
Following previous works [59, 24], we evaluate the effectiveness of our method using empirical robust accuracy (ERA), certified robust accuracy (CRA), average empirical radius (AER), and average certified radius (ACR). Specifically, ERA is defined as the proportion of clean testing samples that are correctly classified by the Cert-SSB model, serving as an upper bound for the corresponding CRA. CRA is defined as the proportion of backdoor-triggered testing samples that are consistently classified correctly within the certified radius $r$. This implies that the trained Cert-SSB model not only ensures that its predictions are consistent with those of a model trained on a clean dataset but also guarantees that these predictions are equal to the ground truth labels. AER represents the average radius over clean test samples, while ACR denotes the average radius over all backdoor-triggered test samples. In general, higher values of ERA, CRA, AER, and ACR indicate better certification performance. In particular, we present certification curves (see Figure 77) to provide a more intuitive comparison of certified performance ($i.e.$, ERA and CRA) under different noise levels.
Figure 6: Certified performance ($i.e.$, ERA and CRA) under different certification radii on the ImagNette dataset in the all-to-one setting with various noise levels (0.25, 0.5, and 1.0). The first column corresponds to the one-pixel attack, the second to the four-pixel attack, and the third to the blending attack.
Figure 7: Certified performance ($i.e.$, ERA and CRA) under different certification radii on the ImagNette dataset in the all-to-all setting with various noise levels (0.25, 0.5, and 1.0). The first column corresponds to the one-pixel attack, the second to the four-pixel attack, and the third to the blending attack.
Figure 8: Effect of stochastic gradient ascent iterations $T$.
Figure 9: Effect of Optimized-Noise Model Count $M_{o}$.
@@ V-B Main Results under the All-to-One Setting
As shown in Tables I-II, our Cert-SSB achieves the best performance under the all-to-one setting across three datasets and three attack types (one-pixel, four-pixel, and blending). For instance, on the MNIST dataset, at a radius of 1.5, ERA exceeds 72% (an improvement of approximately 30%), while CRA surpasses 45% (an increase of around 3%). Even on the more challenging ImageNette dataset, at a radius of 0.75, ERA exceeds 45% (an improvement of nearly 15%), and CRA is above 26% (an increase of 10%). In both cases, AER and ACR also improve by approximately 0.2. These experimental results validate the effectiveness of our certification method.
As shown in Figure 7, our method achieves significantly higher ERA and CRA across different noise levels ($e.g.$, 0.25, 0.5, and 1.0) compared to traditional approaches on the ImageNette dataset, further validating its superior certification performance. Notably, in these scenarios, the trade-off between model accuracy and robustness becomes more pronounced: stronger noise tends to degrade certification performance at smaller radii while improving performance at larger ones. The certification curves for the CIFAR-10 and MNIST datasets are provided in our Appendix -C.
@@ V-C Main Results under the All-to-All Setting
As shown in Tables III-IV, our Cert-SSB method also achieves the best performance under the all-to-all setting across three datasets and three attack types (one-pixel, four-pixel, and blending). For example, on the MNIST dataset, at a radius of 1.5, ERA exceeds 92% (an improvement of approximately 30%), while CRA surpasses 42% (an increase of about 40%). Even on the more challenging ImageNette dataset, at a radius of 0.75, ERA exceeds 40% (an improvement of nearly 15%), and CRA is above 20% (an increase of 20%). In both cases, AER improves by approximately 0.2, while ACR increases by 0.7 on MNIST and 0.4 on ImageNette. These results validate the effectiveness of our method.
As shown in Figure 7, both ERA and CRA under different noise levels ($e.g.$, 0.25, 0.5, 1.0) are also significantly higher than those of traditional methods on the ImageNette dataset, with even more pronounced improvements compared to the all-to-one setting. This experimental result further demonstrates the superior certification performance of our Cert-SSB method. The trade-off between model accuracy and robustness is consistent with that in the all-to-one setting. The certification curves for CIFAR-10 and MNIST are in our Appendix -C.
V-D Discussions
In this section, we discuss the effectiveness of Cert-SSB with different key hyperparameters. For simplicity, we take the one-pixel attack on the ImageNette dataset under the all-to-one setting as an example for discussion.
@@ V-D Discussions
Effect of Optimized-Noise Model Count $\bm{M_{o}}$. Considering that the final prediction is obtained through an ensemble of multiple models, we further investigate the impact of the number of optimized-noise models on certification performance ($i.e.$, ERA and CRA) under different certification radii. Specifically, we trained 50 models with fixed noise $\sigma_{0}$ and 150 models with optimized noise $\sigma_{\bm{x}}^{*}$, forming an ensemble of 200 models ($i.e.$, $M_{f}=50$, $M_{o}=150$). This setup is compared against two baselines: one where all models are trained with fixed noise ($i.e.$, $M_{f}=200$, $M_{o}=0$), and another where all models are trained with optimized noise ($i.e.$, $M_{f}=0$, $M_{o}=200$). We evaluate the certification performance of these three settings under various noise levels ($i.e.$, $\sigma=0.25,0.5,1.0$) and across different certification radii. As shown in Figure 9, the ensemble trained entirely with optimized noise achieves significantly higher ERA and CRA at all certification radii, compared to those incorporating a portion of fixed-noise models. These results indicate that increasing the number of optimized-noise models helps improve the robustness of the ensemble, while introducing fixed-noise models may limit the overall performance.
@@
Visualization of Optimized Noise $\bm{\sigma_{\bm{x}}^{*}}$. We hereby randomly select two input images from two categories, respectively, and perform noise optimization for each input image $\bm{x}$, starting from a fixed initial noise level of $\sigma_{0}=0.25$, to obtain the optimal noise $\sigma_{\bm{x}}^{*}$ that maximizes the certified radius. As shown in Figure 10, the optimized noise values vary significantly across different inputs, with some being larger and others smaller. Notably, even within the same category, there exist considerable differences among the optimized results. These findings further demonstrate the necessity of adaptively optimizing the noise for each individual input.
V-E The Analysis of Computational Complexity

### v2

@@ V Experiments
TABLE III: Certified performance ($i.e.$, ERA and AER) of Cert-SSBD and RAB on MNIST, CIFAR-10, and ImageNette under the all-to-all setting with representative attacks ($i.e.$, one-pixel, four-pixel, and blending). At each radius, we report the best result over different noise levels; the best results are marked in boldface.
@@
Cert-SSBD 0.70 93.32 78.16 42.48 52.34 38.26 17.28 1.40 0
TABLE IV: Certified performance ($i.e.$, CRA and ACR) of Cert-SSBD and RAB on MNIST, CIFAR-10, and ImageNette under the all-to-all setting with representative attacks ($i.e.$, one-pixel, four-pixel, and blending). At each radius, we report the best result over different noise levels; the best results are marked in boldface.
@@ V-A2 Training Settings
We adopt a sample-specific smoothing approach during training. In this stage, we set the number of sampled Gaussian noise vectors ($i.e.$, augmented datasets) to $M=1,000$ for MNIST and CIFAR-10, and $M=200$ for ImageNette, resulting in ensembles of 1,000 and 200 models, respectively. Following previous works [8, 70], the added noise follows a Gaussian distribution with mean $\mu=0$ and a fixed noise level $\sigma_{0}$, set as follows: for MNIST and CIFAR-10, $\sigma_{0}\in\{0.12,0.25,0.5,1.0\}$; for ImageNette, $\sigma_{0}\in\{0.25,0.5,1.0\}$. Additionally, we set the number of stochastic gradient ascent iterations to $T=1$, the number of Monte Carlo samples to $J=1$, and the learning rate to $\alpha=10^{-4}$ (we use $T=100$ during inference unless otherwise specified). During optimization, the sample-specific noise $\sigma_{\bm{x}}^{*}$ is initialized at the fixed noise level $\sigma_{0}$ and is iteratively updated via stochastic gradient ascent.
@@ V-A3 Attack Settings
We evaluate the certified performance of Cert-SSBD against three representative backdoor attacks: one-pixel pattern, four-pixel pattern, and random but fixed noise patterns blended across the entire image [5]. The perturbation magnitude of the attack is controlled by the $\ell_{2}$-norm of the backdoor patterns, with $\left\|\Delta\right\|_{2}=0.1$. Following prior work [59], we inject 10% poisoned samples into the MNIST dataset and 5% into the CIFAR-10 and ImageNette datasets. The goal of these attacks is to induce the model to misclassify inputs as ‘0’ in MNIST, ‘airplane’ in CIFAR-10, and ‘tench’ in ImageNette. In addition to the all-to-one attack, we also consider an all-to-all attack objective [19], where the compromised model alters its predictions based on the original labels. We hereby primarily focus on the perturbation magnitude and the number of injected backdoor samples without considering specific backdoor patterns.
Figure 6: Certified performance ($i.e.$, ERA and CRA) under different certification radii on the ImageNette dataset in the all-to-one setting with various noise levels (0.25, 0.5, and 1.0). The first column corresponds to the one-pixel attack, the second to the four-pixel attack, and the third to the blending attack.
Figure 7: Certified performance ($i.e.$, ERA and CRA) under different certification radii on the ImageNette dataset in the all-to-all setting with various noise levels (0.25, 0.5, and 1.0). The first column corresponds to the one-pixel attack, the second to the four-pixel attack, and the third to the blending attack.
Figure 8: Effect of stochastic gradient ascent iterations $T$.
@@ V-A4 Evaluation Metrics
Following previous works [8, 59], we evaluate the effectiveness of our method using empirical robust accuracy (ERA), certified robust accuracy (CRA), average empirical radius (AER), and average certified radius (ACR). Specifically, ERA refers to the classification accuracy on clean samples (serving as the upper bound for CRA), while CRA denotes the robust accuracy for backdoor samples within the certified radius $r$ ($i.e.$, the predictions are provably invariant within $r$ and correct). AER is the average empirical radius over clean samples, while ACR is the average certified radius over backdoor samples. In general, higher values of ERA, CRA, AER, and ACR indicate better certification performance. In particular, we present certification curves (see Figure 77) to intuitively compare certified performance ($i.e.$, ERA and CRA) under different noise levels.
@@ V-B Main Results under the All-to-One Setting
As shown in Tables I-II, our Cert-SSBD achieves the best performance under the all-to-one setting across three datasets and three attack types (one-pixel, four-pixel, and blending). For instance, on the MNIST dataset, at a radius of 1.5, ERA exceeds 72% (an improvement of approximately 30%), while CRA surpasses 45% (an increase of around 3%). Even on the more challenging ImageNette dataset, at a radius of 0.75, ERA exceeds 45% (an improvement of nearly 15%), and CRA is above 26% (an increase of 10%). In both cases, AER and ACR also improve by approximately 0.2. These experimental results validate the effectiveness of our certification method.
As shown in Figure 7, our method achieves significantly higher ERA and CRA across various noise levels ($e.g.$, 0.25, 0.5, and 1.0) on ImageNette compared to traditional methods, validating its superior performance. Notably, the trade-off between accuracy and robustness is more pronounced in this context: stronger noise tends to degrade performance at smaller radii while improving it at larger ones. The certification curves for CIFAR-10 and MNIST are provided in Appendix -D.
@@ V-C Main Results under the All-to-All Setting
As shown in Tables III-IV, our Cert-SSBD method also achieves the best performance under the all-to-all setting across three datasets and three attack types (one-pixel, four-pixel, and blending). For example, on the MNIST dataset, at a radius of 1.5, ERA exceeds 92% (an improvement of approximately 30%), while CRA surpasses 42% (an increase of about 40%). Even on the more challenging ImageNette dataset, at a radius of 0.75, ERA exceeds 40% (an improvement of nearly 15%), and CRA is above 20% (an increase of 20%). In both cases, AER improves by approximately 0.2, while ACR increases by 0.7 on MNIST and 0.4 on ImageNette. These results validate the effectiveness of our method.
As shown in Figure 7, our method achieves significantly higher ERA and CRA on ImageNette under various noise levels ($e.g.$, 0.25, 0.5, and 1.0) compared to traditional methods. Furthermore, the trade-off between model accuracy and robustness remains consistent with the all-to-one setting. The curves for CIFAR-10 and MNIST are provided in Appendix -D.
TABLE V: Certified performance ($i.e.$, ERA and AER) of Cert-SSBD and RAB on MNIST under the all-to-one setting with diverse trigger designs ($i.e.$, BadNets, WaNet, SIG, and an adaptive trigger). At each radius, we report the best result over different noise levels; the best results are marked in boldface.
[…]
0 0.25 0.5 0.75 1.0 1.25 1.5 1.75
BadNets RAB 1.56 99.95 99.86 99.76 99.67 99.43 98.82 74.80 0
Cert-SSBD 1.81 100 100 100 100 100 99.57 99.15 98.49
WaNet RAB 1.49 100 99.91 99.67 99.57 98.96 50.83 0 0
Cert-SSBD 1.56 100 99.86 99.81 99.34 98.91 97.07 78.16 39.67
SIG RAB 1.54 99.95 99.86 99.76 99.67 99.48 93.48 0 0
Cert-SSBD 1.74 99.95 99.86 99.86 99.67 99.34 98.82 96.97 88.09
Adaptive Trigger RAB 1.51 99.95 99.86 99.76 99.57 99.20 98.44 79.43 0
Cert-SSBD 1.65 100 99.86 99.81 99.62 99.20 98.53 93.52 41.70
TABLE VI: Certified performance ($i.e.$, CRA and ACR) of Cert-SSBD and RAB on MNIST under the all-to-one setting with diverse trigger designs ($i.e.$, BadNets, WaNet, SIG, and an adaptive trigger). At each radius, we report the best result over different noise levels; the best results are marked in boldface.
[…]
0 0.25 0.5 0.75 1.0 1.25 1.5 1.75
BadNets RAB 1.49 99.95 99.86 99.72 99.48 99.01 94.52 0 0
Cert-SSBD 1.74 99.95 99.86 99.86 99.62 99.34 98.77 96.88 87.09
WaNet RAB 0.68 91.96 46.24 46.01 46.01 45.86 45.25 39.72 0
Cert-SSBD 0.79 46.29 46.24 46.15 45.91 45.82 45.01 43.22 37.92
SIG RAB 1.53 99.95 99.91 99.86 99.72 99.34 98.72 90.26 0
Cert-SSBD 1.73 99.95 99.86 99.86 99.72 99.43 98.82 96.64 86.95
Adaptive Trigger RAB 0.68 46.76 46.29 46.01 46.01 45.91 45.58 41.89 0
Cert-SSBD 0.81 99.81 97.26 46.34 46.01 45.91 45.67 44.44 41.23
V-D Ablation Study
In this section, we conduct ablation studies to analyze the effects of key design choices in Cert-SSBD. Unless otherwise specified, experiments are conducted using the one-pixel attack on the ImageNette dataset under the all-to-one setting. More ablation results are provided in Appendix -H.
@@ V-D Ablation Study
Effect of Trigger Diversity on Certified Robustness. We hereby evaluate Cert-SSBD under a more diverse set of backdoor trigger settings, including BadNets [19], WaNet [47], SIG [3], and an adaptive trigger [48]. We hereby use the all-to-one setting on MNIST as an example for discussion. All other training and certification settings follow Section V-A to ensure fair comparisons. Specifically, we adopt a fixed patch trigger placed at the bottom-right corner for BadNets; use the smooth geometric warping-based trigger of WaNet; inject a globally diffused low-amplitude sinusoidal signal as in SIG; and employ an input-aware adaptive trigger whose pattern and placement are jointly optimized with respect to the target model. As shown in Tables V-VI, Cert-SSBD consistently outperforms RAB across all four trigger designs, with more pronounced advantages at larger certification radii. Taking BadNets as an example, at radius $r=1.5$, Cert-SSBD achieves an ERA of 99.15% with a corresponding CRA of 96.88%, whereas RAB attains an ERA of 74.80% and its CRA drops to 0. Meanwhile, the AER increases from 1.56 to 1.81, and the ACR improves from 1.49 to 1.74. For the adaptive trigger, at a larger radius $r=1.75$, both the ERA and CRA of RAB drop to 0, while Cert-SSBD still maintains an ERA of 41.70% and a CRA of 41.23%. In addition, the AER improves from 1.51 to 1.65, and the ACR increases from 0.68 to 0.81. These results demonstrate that Cert-SSBD remains robust and consistent under more diverse and challenging trigger settings, especially at larger certification radii.
@@
Figure 10: Distribution of optimized noise $\sigma_{x}^{*}$ on the MNIST testing set under different fixed noise levels $\sigma_{0}$ for clean and poisoned testing samples. The left panel shows clean testing samples, while the right panel shows poisoned testing samples under the one-pixel attack.
V-E Discussions
In this section, we provide further analysis and discussion to better understand the behavior, interpretability, and robustness of the optimized noise $\sigma_{\bm{x}}^{*}$ under diverse conditions.
V-E1 Visualization of Optimized Noise
@@
We hereby randomly select two input images from two categories, respectively, and perform noise optimization for each input image $\bm{x}$, starting from a fixed initialized noise level of $\sigma_{0}=0.25$, to obtain the optimal noise $\sigma_{\bm{x}}^{*}$ that maximizes the certified radius. As shown in Figure 9, the optimized noise values vary significantly across different inputs, with some being larger and others smaller. Notably, even within the same category, there exist considerable differences among the optimized results. These findings further demonstrate the necessity of adaptively optimizing the noise for each individual input.
V-E2 Distribution of Optimized Noise
We hereby analyze the distribution of the optimized per-sample noise $\sigma_{\bm{x}}^{*}$, considering both clean test samples and poisoned test samples generated by a one-pixel attack. For discussion, we use the MNIST dataset under the all-to-one setting as an illustrative example. We report the results under different fixed noise levels, with $\sigma_{0}\in\{0.12,0.25,0.5,1.0\}$. As shown in Figure 10, the optimized noise exhibits a non-uniform distribution across the dataset under all noise levels, indicating that different samples are assigned distinct noise magnitudes. Moreover, as the fixed noise level $\sigma_{0}$ increases, the overall range of $\sigma_{\bm{x}}^{*}$ correspondingly expands. Nevertheless, even under the same fixed noise level, substantial variability persists among individual samples. Therefore, the proposed optimization procedure adaptively adjusts $\sigma_{\bm{x}_{i}}^{*}$ for each sample, rather than learning a single fixed noise level.
TABLE VII: Certified performance ($i.e.$, ERA and AER) of Cert-SSBD on MNIST under the all-to-one setting against margin-aware adaptive poisoning (MAP) combined with representative attacks ($i.e.$, one-pixel, four-pixel, and blending). At each radius, we report the best result over different noise levels; the best certification results are marked in boldface.
[…]
0 0.25 0.5 0.75 1.0 1.25 1.5 1.75 2.0
One-pixel 1.65 99.95 99.91 99.81 99.62 99.34 98.82 86.53 42.98 0
One-pixel + MAP 1.78 99.91 99.86 99.86 99.53 98.63 95.70 85.39 61.23 24.21
Four-pixel 1.69 99.95 99.86 99.72 99.57 99.20 98.63 81.94 42.98 0
Four-pixel + MAP 1.75 99.91 99.86 99.81 99.39 97.97 94.14 82.08 56.69 23.64
Blending 1.70 99.95 99.86 99.76 99.72 99.20 98.72 72.15 42.84 0
Blending + MAP 1.83 99.91 99.91 99.86 99.62 98.06 94.47 84.54 65.96 36.93
TABLE VIII: Certified performance ($i.e.$, CRA and ACR) of Cert-SSBD on MNIST under the all-to-one setting against margin-aware adaptive poisoning (MAP) combined with representative attacks ($i.e.$, one-pixel, four-pixel, and blending). At each radius, we report the best result over different noise levels; the best certification results are marked in boldface.
[…]
0 0.25 0.5 0.75 1.0 1.25 1.5 1.75 2.0
One-pixel 0.84 46.29 46.24 46.24 46.10 45.96 45.91 45.20 42.88 0
One-pixel + MAP 0.90 90.00 90.40 47.16 46.15 45.91 45.77 44.78 40.80 24.16
Four-pixel 0.87 46.29 46.24 46.24 46.10 46.01 45.91 45.67 43.88 0
Four-pixel + MAP 0.89 93.99 92.01 85.20 46.24 45.86 45.62 44.21 38.91 23.12
Blending 0.87 46.34 46.29 46.24 46.19 45.96 45.91 45.63 44.30 0
Blending + MAP 0.95 92.67 87.85 47.47 46.19 45.91 45.82 45.20 43.36 36.40
V-E3 Result to Potential Adaptive Attack on Optimized Noise
Assuming an adaptive adversary who is fully aware of our defense mechanism, the attacker may attempt to weaken the certified robustness of Cert-SSBD in an indirect manner. Specifically, in Cert-SSBD, the noise scale $\sigma_{\bm{x}}$ for each sample $\bm{x}_{i}$ is optimized via stochastic gradient ascent (SGA), and its optimal value $\sigma_{\bm{x}_{i}}^{*}$ is closely related to the sample’s proximity to the model’s decision boundary. Therefore, an adaptive attacker can perform targeted poisoning of the training data to push the decision boundary closer to selected target samples, thereby indirectly reducing the statistical advantage ($e.g.$, $P_{A}-P_{B}$) and ultimately shrinking the certified radius. To evaluate the robustness of Cert-SSBD under such adaptive adversary scenarios, we design a Margin-Aware Adaptive Poisoning (MAP) attack to simulate this capability. Unlike standard poisoning attacks that randomly select training samples, MAP strategically selects poisoning samples based on their proximity to a set of vulnerable testing samples, where vulnerability is identified via the logit margin. It is worth emphasizing that MAP uses the same trigger patterns as standard attacks ($e.g.$, one-pixel, four-pixel, and blending); its key distinction lies solely in the poisoning sample selection strategy rather than the trigger pattern itself. Specifically, MAP adopts a two-stage heuristic strategy: (1) identifying vulnerable testing samples with small logit margins ($i.e.$, close to the decision boundary), and (2) selecting training samples with the smallest feature distances to these vulnerable samples for poisoning. The detailed formalization and implementation of the MAP attack are provided in Appendix -G.
As shown in Tables VII-VIII, Cert-SSBD demonstrates good robustness against the MAP adaptive attack. The adaptive poisoning strategy does not significantly degrade the certification performance, and some metrics even show improvements. On the clean testing set (Table VII), ERA exhibits slight fluctuations at intermediate radii but remains overall stable, with a notable improvement at $r=2.0$, increasing from 0% to 24.21%–36.93%. AER improves to varying degrees under all attacks, with blending + MAP reaching 1.83 (compared to 1.70 for standard poisoning). On the poisoned testing set (Table VIII), changes are more pronounced at small and large radii. At $r=0$, CRA shows substantial growth ($e.g.$, one-pixel + MAP increases from 46.29% to 90.00%); at $r=2.0$, CRA improves from 0% to 23.12%–36.40%. ACR improves overall, with blending + MAP reaching 0.95 (compared to 0.87 for standard poisoning). These results indicate that the sample-specific noise optimization mechanism of Cert-SSBD exhibits a degree of inherent robustness: even when adaptive poisoning indirectly perturbs the model parameters, the SGA-based optimization of $\sigma_{\bm{x}}^{*}$ can still adaptively adjust the noise scale, thereby empirically maintaining relatively stable certified defense performance.
V-F The Analysis of Computational Complexity